# Panoramic Out-of-Distribution Segmentation for Autonomous Driving

### v2

@@ III-D Bilevel Prompt Distribution Learning
As demonstrated in prior works [78, 79], increasing the inter-class distance between category centers helps mitigate semantic ambiguity across different classes. Building upon this insight, similarly, an inter-class separation loss $\mathcal{L}_{inter}$ is proposed to refine the classification boundaries in the feature space. Unlike existing approaches that rely on learned or manually defined class centers, our method leverages semantically clear text prompts as predefined category prototypes, which inherently exhibit better separation due to their linguistic distinctness. The proposed $\mathcal{L}_{inter}$ then acts as a regularization term to further constrain the feature distribution, ensuring structured separation and preventing uncontrolled dispersion of embeddings. We primarily propose an separation loss $\mathcal{L}_{sep}$ that increases the distance between each class center $T_{i}$ and its nearest negative class center $T_{n}$:
$\mathcal{L}_{sep}=\frac{1}{K+1}\sum_{i=0}^{K}\max(s-\min_{n\neq i}\left\|T_{i}-T_{n}\right\|_{2}^{2},~0).$ (9)
The distance margin $s$ is defined as a hyperparameter. In addition, to account for the spatial relationships among the remaining negative class centers in the feature space, we introduce an additional orientation constraint $\mathcal{L}_{ori}$ to optimize the directional relationships of class embedding distributions. As shown in Fig. 7 (right), when repelling the nearest negative class center $T_{n}$, the angle $\theta$ between the direction $\vec{t}_{n}$ (from $T_{i}$ to $T_{n}$) and the direction $\vec{t}_{o}$ (from $T_{i}$ to other negative class centers $T_{o}$) tends to increase:
$\vec{t}_{n}=T_{n}-T_{i},$ (10)
$\vec{t}_{o}=T_{o}-T_{i},$
$\mathcal{L}_{ori}=\frac{1}{K+1}\sum_{i=0}^{K}\sum_{o\neq i,n}\omega_{o}(1+\cos(\vec{t}_{n},~\vec{t}_{o})),$ (11)
where $\omega_{o}$ represents the weights assigned to other negative class centers $T_{o\neq i,n}$, which are determined based on their spatial relationships relative to $T_{n}$:
$\omega_{o}=\frac{\exp(-\|T_{n}-T_{o}\|_{2})}{\sum_{k\neq i,n}\exp(-\|T_{n}-T_{k}\|_{2})}.$ (12)
The inter-class separation loss $\mathcal{L}_{inter}$ is formulated as combination of two key terms:
@@
$\mathcal{L}_{pixel}=\mathcal{L}_{intra}+\mathcal{L}_{inter},$ (14)
The $\mathcal{L}_{pixel}$ facilitates that, while capturing the overall feature space distribution, the model further refines its ability to differentiate in-distribution class distributions and enhances its semantic understanding.
Distribution learning. Building upon the aforementioned learning, a higher-level distribution learning mechanism is proposed to enhance the separability of out-of-distribution embeddings $O_{i}$ and achieve a more precise delineation of the overall feature space. To establish a robust foundation for semantic alignment in subsequent distribution region partitioning and enable the model to more accurately capture the feature distribution patterns of known classes, we first minimize the distance between class text prompts $\{T_{i}\mid i=0,\dots,K\}$ and in-distribution prompts $P_{in}$ in the feature space. The in-distribution compactness loss $\mathcal{L}_{ind}$ promotes that the feature distribution of known classes aligns with the in-distribution semantic information:
$\mathcal{L}_{ind}=\frac{1}{K+1}\sum_{i=0}^{K}\left\|T_{i}-P_{in}\right\|_{2}^{2}.$ (15)
To further constrain the distributional difference between known class features and unknown pixel features, we propose a directional constraint strategy. As depicted in Fig. 8, our goal is to position known class embeddings in the opposite direction of the out-of-distribution prompt $P_{out}$. Specifically, we define the direction from known class embeddings to $P_{in}$ as $\vec{d}_{in}$ and the direction to $P_{out}$ as $\vec{d}_{out}$. We optimize these directions to be as aligned as possible by maximizing their cosine similarity. The $\vec{d}_{in}$ and $\vec{d}_{out}$ are formulated as:
$\vec{d}_{in}$ $=P_{in}-T_{i},$ (16)
$\vec{d}_{out}$ $=P_{out}-T_{i}.$
This approach effectively mitigates confusion between in-distribution and out-of-distribution pixel embeddings in the feature space, thereby enhancing the model’s robustness to unknown features. The in-distribution directional loss $\mathcal{L}_{inDir}$ is defined as follows:
$\mathcal{L}_{inDir}=\frac{1}{K+1}\sum_{i=0}^{K}1-\cos(\vec{d}_{in},~\vec{d}_{out}).$ (17)
The overall in-distribution loss $\mathcal{L}_{inlier}$ is formulated as the weighted combination of the in-distribution compactness loss $\mathcal{L}_{ind}$ and directional loss $\mathcal{L}_{inDir}$:
$\mathcal{L}_{inlier}=\alpha~\mathcal{L}_{ind}+\mathcal{L}_{inDir},$ (18)
@@
Figure 8: Distribution Learning. Isolation of out-of-distribution pixels: A distinct boundary is established in the feature space to separate out-of-distribution regions from known class regions, thereby reducing the overconfidence for outliers. Aggregation of out-of-distribution pixels: through semantic guidance, out-of-distribution pixel embeddings are clustered into specific semantic spaces, preventing their dispersion and interference with the feature distribution of known classes.
Furthermore, to enhance the model’s ability to identify out-of-distribution regions, we employ semantic guidance to cluster out-of-distribution pixel embeddings $\{O_{1},\dots,O_{N_{o}}\}$ into specific regions of the semantic space. This approach not only preserves the stability of the known class feature distribution but also provides clear semantic attribution for out-of-distribution features. Additionally, it distinctly delineates out-of-distribution regions in the feature space, establishing a clear boundary with known class regions. As a result, it reduces the model’s overconfidence for outliers and enhances the reliability of detection.
During the fine-tuning, outliers are pasted into the input images with a preset probability. This outlier exposure strategy simulates the appearance of out-of-distribution samples in real-world scenarios, enabling the model to better adapt to complex tasks in open environments. To facilitate model learning, we propose an out-of-distribution pixel loss $\mathcal{L}_{outlier}$ for both the presence and absence of outliers:
@@
where $\{D_{i,out},~D_{i,in}\}=\{\|O_{i}-P_{out}\|_{2}^{2},~\|O_{i}-P_{in}\|_{2}^{2}\}$, $d$ quantifies the degree of difference between $P_{in}$ and $P_{out}$ within the semantic space. The distribution learning loss $\mathcal{L}_{distri}$ is formulated as a sum of the in-distribution loss $\mathcal{L}_{inlier}$ and the out-of-distribution pixel loss $\mathcal{L}_{outlier}$, which is calculated as follows:
@@
Finally, the bilevel prompt distribution learning loss is defined as follows:
@@
Through the bilevel prompt distribution learning mechanism, the manifold of per-pixel embeddings for mask segmentation becomes more compact and distinguishable, thereby demonstrating better robustness and generalization in PanOoS.
@@ III-E Loss Settings
Following [29, 27, 28], during training, bipartite matching is employed to match each ground truth mask with an object query, supervised by region prediction and classification losses, while the fine-tuning phase further incorporates outlier supervision and bilevel prompt distribution learning. For region prediction, a weighted combination of binary cross-entropy loss and Dice loss [80]:
@@
In the outlier exposure fine-tuning phase, we further utilize the $RbA$ loss $\mathcal{L}_{rba}$ [27] to suppress the high-confidence probabilities of outlier pixels in known classes, and the bilevel prompt distribution learning loss $\mathcal{L}_{bpdl}$ to refine the pixel manifold of per-pixel mask embeddings:
@@
where $\lambda_{bdpl}$ is the balancing coefficient.
@@ IV-A Overview of the Benchmarks
In this work, we establish two novel benchmarks: DenseOoS and QuadOoS, specifically designed for PanOoS. $360^{\circ}$ panoramas have a broader FoV and many small objects in the images, which exaggerate the costs of creating pixel-wise annotations in unconstrained surroundings. From the perspective of the task, we aim to overcome the scarcity of PanOoS testbeds and effectively evaluate the impact of mitigating the background clutter and severe distortions on out-of-distribution segmentation performance when unmasking the narrow FoV. Furthermore, to explore PanOoS in real-world, unconstrained, unstructured, and dense-motion scenarios, we utilize a quadruped mobile robot to capture images in outdoor motion scenes with outliers, creating a real-world benchmark. Our objective with two benchmarks is to provide a comprehensive evaluation of methods capable of performing PanOoS. For comparison to previous pinhole-OoS methods, we adapt Area under PRC Curve (AuPRC/AP), and False Positive Rate at a true positive rate of $95\%$ (FPR95) for panoramic out-of-distribution segmentation pixel-wise evaluation, mean Intersection-over-Union (mIoU) for closed-set segmentation evaluation. We unfold a comprehensive explanation in Sec V-C.
@@
Road Anomaly [51] 2019 Pinhole outlier
Fishscapes L&F [52] 2019 outlier (✓)
Fishscapes Static [52] 2019 outlier
StreetHazards [55] 2019 all (✓)
SMIYC RA-21 [50] 2021 outlier
SMIYC RO-21 [50] 2021 outlier
@@
TABLE I: Comparison of different OoS datasets. Referring to POC [81], we qualitatively compare datasets on two main axes, and score them as good (), medium ((✓)) or bad ().
@@ IV-B DenseOoS
We introduce a novel test dataset: Dense Panoramic Out-of-Distribution Segmentation (DenseOoS) tailored for PanOoS. DenseOoS comprises a labeled test set of $1,000$ panoramic images with a size of $2048{\times}400$ pixels and provides pixel-wise annotations for the PanOoS task, which greatly extends semantic labels from DensePASS [18]. Relevant studies have shown that using generated outlier data for model training or evaluation is an effective approach [82, 81]. Therefore, based on the DensePASS dataset, we utilize a state-of-the-art image generation technique: POC [81], to place multiple classes of outliers in images while generating annotation labels for these outliers to construct DenseOoS.
The normal labels cover $19$ classes that align with the Cityscapes [15], while the abnormal labels contain $30$ classes of outliers as demonstrated in Fig. 9. To ensure the high quality and rationality of the generated images, particularly that the outliers are generated in realistic areas within the real street backgrounds, our screening process follows the “preliminary screening $\rightarrow$} independent screening $\rightarrow$} voting” workflow, eliminating unreasonable object locations, and normal situations. Specifically, a total of $14,031$ images are generated, of which $1,244$ are retained after preliminary manual screening. Following an independent screening and voting process by three examiners, we ultimately select $1,000$ images for the construction of DenseOoS. More details are in the Appendix. As shown in Fig. 10, the outliers are diverse and distributed within reasonable areas of the images. A significant portion of these outliers extends beyond the pinhole narrow FoV, making them particularly suitable for the evaluation of the PanOoS task.
@@ IV-C QuadOoS
To further explore the challenges of the PanOoS task in real applications, we collect panoramic images featuring outliers in unconstrained, unstructured, and dense-motion real scenes, manually annotating these outliers to establish the first Quadruped Panoramic Out-of-distribution Segmentation dataset, QuadOoS. This dataset provides a crucial foundation for panoramic out-of-distribution segmentation research, facilitating the evaluation and optimization of PanOoS models in complex environments. We develop a mobile platform using a quadruped robot, and the robot was selected for its biomimetic gait, a complex form of locomotion found in nature, which introduces additional challenges for PanOoS. It can navigate vertical obstacles up to $15cm$ and inclines up to $30^{\circ}$, making it highly maneuverable in everyday environments. With $12$ joint motors, the robot replicates realistic walking motions at speeds up to $2.5m/s$. For sensing, we use a customized panoramic annular camera to capture a broader FoV ($360^{\circ}{\times}70^{\circ}$). The camera has a pixel size of $3.45{\mu}m{\times}3.45{\mu}m$, a resolution of $5$ million effective pixels. Mounted on the quadruped robot (see Fig. 11(a)), the camera ensures an optimal field of view.
Figure 10: Examples from the DenseOoS dataset. The background of the images is highly cluttered, and the distortions are significant. Outliers are distributed in reasonable areas within the image. Zoom in for a better view.
@@
Using this platform, we conduct data collection in campus environments with outliers, creating a Quadruped Panoramic Out-of-distribution Segmentation (QuadOoS) dataset. In particular, we collect data across $8$ blocks in and around the campus, with each block averaging approximately $50m$ in length. The collected images consist of a total of $4,800$ frames ($600$ frames per block), each with an original size of $2048{\times}476$ pixels. To avoid misleading the model due to excessive exposure at the center of the lens and to maintain consistency with DenseOoS, we crop the upper portion of the original images (i.e., the sky), which does not affect fair evaluation. Subsequently, to ensure the diversity of outliers and the satisfactory imaging quality, we retain a final selection of $106$ original panoramic images after manual screening by five examiners, thereby constructing the QuadOoS dataset. Detailed information is provided in Fig. 11(b), and see Appendix B-A for details of the types of outliers.
@@
Due to the gait of the quadruped robot, the collected panoramic images exhibit irregular shaking, particularly along the Y-axis (see Fig. 11(c)). Compared to panoramic images captured under static conditions, the images in QuadOoS present significant challenges to existing methods, such as uneven exposure, color inconsistencies due to the broader FoV, and increased motion blur, as rapid relative displacement between moving objects and the background intensifies the blurring effect. As shown in Fig. 12, each image contains an average of $7$ outliers, accurately reflecting real-world situations and posing substantial challenges for PanOoS methods. As illustrated in Table I, with a broader FoV ($360^{\circ}{\times}70^{\circ}$), QuadOoS significantly differs from traditional pinhole-OoS datasets [51, 52, 55, 50], and complements DenseOoS by enabling the exploration of motion-induced and panoramic-specific challenges in PanOoS.
@@
Figure 12: Examples from the established QuadOoS dataset. The images come with challenges such as uneven exposure, color inconsistencies, and increased motion blur. A large number of small outliers are scattered in the background.

### v3

@@ III-D1 In-distribution learning
As demonstrated in prior works [he2018triplet, liu2022orientational], increasing the inter-class distance between category centers helps mitigate semantic ambiguity. Building upon this insight, we propose an inter-class separation loss $\mathcal{L}_{inter}$ to refine the classification boundaries in the feature space. Unlike existing approaches that rely on learned or manually defined class centers, our method leverages text prompts as predefined category prototypes, which inherently exhibit better separation due to their linguistic distinctness. We primarily propose a separation loss $\mathcal{L}_{sep}$ that increases the distance between each class center $\mathbf{T}_{i}$ and its nearest negative class center $\mathbf{T}_{n}$:
$\mathcal{L}_{sep}=\frac{1}{K+1}\sum_{i=0}^{K}\max(s-\min_{n\neq i}\left\|\mathbf{T}_{i}-\mathbf{T}_{n}\right\|_{2}^{2},\,0).$ (10)
The distance margin $s$ is defined as a hyperparameter.
In addition, to account for the spatial relationships among the remaining negative class centers in the feature space, we introduce an orientation constraint $\mathcal{L}_{ori}$ to optimize the directional relationships of class embedding distributions. As illustrated in Fig. 7 (right), when repelling the nearest negative class center $\mathbf{T}_{n}$, the angle $\theta$ between the direction $\vec{t}_{n}$ (from $\mathbf{T}_{i}$ to $\mathbf{T}_{n}$) and the direction $\vec{t}_{o}$ (from $\mathbf{T}_{i}$ to other negative class centers $\mathbf{T}_{o}$) tends to increase:
$\vec{t}_{n}=\mathbf{T}_{n}-\mathbf{T}_{i},$ (11)
$\vec{t}_{o}=\mathbf{T}_{o}-\mathbf{T}_{i},$
$\mathcal{L}_{ori}=\frac{1}{K+1}\sum_{i=0}^{K}\sum_{o\neq i,n}\omega_{o}\left(1+\cos\left(\vec{t}_{n},\vec{t}_{o}\right)\right),$ (12)
where $\omega_{o}$ represents the weights assigned to other negative class centers $\mathbf{T}_{o}$ (with $o\neq i,n$), which are determined based on their spatial relationships relative to $\mathbf{T}_{n}$:
$\omega_{o}=\frac{\exp\left(-\left\|\mathbf{T}_{n}-\mathbf{T}_{o}\right\|_{2}\right)}{\sum_{k\neq i,n}\exp\left(-\left\|\mathbf{T}_{n}-\mathbf{T}_{k}\right\|_{2}\right)}.$ (13)
The inter-class separation loss $\mathcal{L}_{inter}$ is formulated as a combination of two key terms:
@@
This acts as a regularization term to further constrain the feature distribution, ensuring structured separation and preventing uncontrolled dispersion of embeddings.
@@
$\mathcal{L}_{pixel}=\mathcal{L}_{intra}+\mathcal{L}_{inter}.$ (15)
Minimizing the loss $\mathcal{L}_{pixel}$ enables the model, while capturing the overall feature space distribution, to further refine its ability to differentiate between in-distribution class distributions.
Figure 8: Distribution Learning. Isolation of out-of-distribution pixels: A distinct boundary is established in the feature space to separate out-of-distribution regions from known class regions, thereby mitigating overconfident predictions on outliers. Aggregation of out-of-distribution pixels: Through semantic guidance, out-of-distribution pixel embeddings are clustered into specific semantic spaces, preventing their dispersion and interference with the feature distribution of known classes.
III-D2 Distribution learning
Building upon the aforementioned learning, a higher-level distribution learning mechanism is proposed to enhance the separability of out-of-distribution embeddings $\mathbf{O}_{i}$ and achieve a more precise delineation of the overall feature space. To establish a robust foundation for semantic alignment in subsequent distribution region partitioning and enable the model to more accurately capture the feature distribution patterns of known classes, we first minimize the distance between class text prompts $\{\mathbf{T}_{i}\mid i=0,\dots,K\}$ and in-distribution prompt $\mathbf{P}_{in}$ in the feature space:
$\mathcal{L}_{ind}=\frac{1}{K+1}\sum_{i=0}^{K}\left\|\mathbf{T}_{i}-\mathbf{P}_{in}\right\|_{2}^{2}.$ (16)
The in-distribution compactness loss $\mathcal{L}_{ind}$ promotes alignment between the feature distribution of known classes and the in-distribution semantic information.
To further constrain the distributional discrepancy between known class features and unknown pixel features, we propose a directional constraint strategy. As depicted in Fig. 8, our goal is to position known class embeddings in the opposite direction of the out-of-distribution prompt $\mathbf{P}_{out}$. Specifically, we define the direction from known class embeddings to $\mathbf{P}_{in}$ as $\vec{d}_{in}$ and the direction to $\mathbf{P}_{out}$ as $\vec{d}_{out}$. We optimize these directions to be as aligned as possible by maximizing their cosine similarity. The $\vec{d}_{in}$ and $\vec{d}_{out}$ are formulated as:
$\vec{d}_{in}$ $=\mathbf{P}_{in}-\mathbf{T}_{i},$ (17)
$\vec{d}_{out}$ $=\mathbf{P}_{out}-\mathbf{T}_{i}.$
The in-distribution directional loss $\mathcal{L}_{inDir}$ is defined as:
$\mathcal{L}_{inDir}=\frac{1}{K+1}\sum_{i=0}^{K}\left(1-\cos\left(\vec{d}_{in},\vec{d}_{out}\right)\right).$ (18)
This approach effectively mitigates confusion between in-distribution and out-of-distribution pixel embeddings in the feature space, thereby enhancing the model’s robustness to unknown features. The overall in-distribution loss $\mathcal{L}_{inlier}$ is formulated as the weighted combination of the in-distribution compactness loss $\mathcal{L}_{ind}$ and directional loss $\mathcal{L}_{inDir}$:
$\mathcal{L}_{inlier}=\alpha\,\mathcal{L}_{ind}+\mathcal{L}_{inDir},$ (19)
@@ III-D2 Distribution learning
Furthermore, to enhance the model’s ability to identify out-of-distribution regions, we employ semantic guidance to cluster out-of-distribution pixel embeddings $\{\mathbf{O}_{1},\dots,\mathbf{O}_{N_{o}}\}$ into specific regions of the semantic space. During the fine-tuning, outliers are pasted into the input images with a preset probability. This outlier exposure strategy simulates the appearance of out-of-distribution samples in real-world scenarios, enabling the model to better adapt to complex tasks in open environments. To facilitate model learning, we propose an out-of-distribution pixel loss $\mathcal{L}_{outlier}$ for both the presence and absence of outliers:
@@
where $\{\mathbf{D}_{i,out},~\mathbf{D}_{i,in}\}=\{\|\mathbf{O}_{i}-\mathbf{P}_{out}\|_{2}^{2},~\|\mathbf{O}_{i}-\mathbf{P}_{in}\|^{2}_{2}\}$, $d$ quantifies the degree of difference between $\mathbf{P}_{in}$ and $\mathbf{P}_{out}$ within the semantic space. This approach not only preserves the stability of the known class feature distribution but also provides clear semantic attribution for out-of-distribution features. Additionally, it distinctly delineates out-of-distribution regions in the feature space, establishing a clear boundary with known class regions. As a result, it reduces the model’s overconfidence toward outliers and enhances the reliability of detection. The distribution learning loss $\mathcal{L}_{distri}$ is formulated as a sum of the in-distribution loss $\mathcal{L}_{inlier}$ and the out-of-distribution pixel loss $\mathcal{L}_{outlier}$:
@@
Finally, the BPDL loss $\mathcal{L}_{bpdl}$ is defined as follows:
@@
Through the Bilevel Prompt Distribution Learning mechanism, the manifold of per-pixel embeddings for mask predictions becomes more compact and structured, thereby demonstrating better robustness and generalization in PanOoS.
@@ III-E Loss Settings
Following [cheng2022masked, nayal2023rba, rai2024mask2anomaly], during training, bipartite matching is employed to match each ground truth mask with an object query, supervised by region prediction and classification losses, while the fine-tuning phase further incorporates outlier supervision and bilevel prompt distribution learning. For region prediction, a weighted combination of binary cross-entropy loss and Dice loss [milletari2016v]:
@@
In the outlier exposure fine-tuning phase, we further utilize the $RbA$ loss $\mathcal{L}_{rba}$ [nayal2023rba] to suppress the high-confidence probabilities of outlier pixels in known classes, and the bilevel prompt distribution learning loss $\mathcal{L}_{bpdl}$ to refine the pixel manifold of per-pixel mask embeddings:
@@
where $\lambda_{bpdl}$ is the balancing coefficient.
@@ IV-A Overview of the Benchmarks
In this work, we establish two novel benchmarks: DenseOoS and QuadOoS, specifically designed for PanOoS. $360^{\circ}$ panoramas have a broader FoV and many small objects, which exaggerate the costs of creating pixel-wise annotations in unconstrained surroundings. From the perspective of the task, we aim to overcome the scarcity of PanOoS testbeds and effectively evaluate the impact of mitigating the severe distortions and background clutter on out-of-distribution segmentation performance when unmasking the narrow FoV. Furthermore, to explore PanOoS in real-world, unconstrained, unstructured, and dense-motion scenarios, we deploy a quadruped mobile robot to capture images in outdoor motion scenes with outliers, creating a real-world benchmark. Our objective with two benchmarks is to provide a comprehensive evaluation of methods capable of performing PanOoS. For comparison with previous pinhole-OoS methods, we adopt Area under PRC Curve (AuPRC/AP) and False Positive Rate at a true positive rate of $95\%$ (FPR95) for panoramic out-of-distribution segmentation pixel-wise evaluation, and mean Intersection-over-Union (mIoU) for closed-set segmentation evaluation. We unfold a comprehensive explanation in Appendix B-B.
@@
Road Anomaly [lis2019detecting] 2019 Pinhole outlier
Fishscapes L&F [blum2021fishyscapes] 2019 outlier (✓)
Fishscapes Static [blum2021fishyscapes] 2019 outlier
StreetHazards [hendrycks2019scaling] 2019 all (✓)
SMIYC RA-21 [chan2021segmentmeifyoucan] 2021 outlier
SMIYC RO-21 [chan2021segmentmeifyoucan] 2021 outlier
@@
TABLE I: Comparison of different OoS datasets. Referring to POC [de2024placing], we qualitatively compare datasets on two main axes, and score them as good (), medium ((✓)) or bad ().
@@ IV-B DenseOoS
We introduce a novel test dataset: Dense Panoramic Out-of-Distribution Segmentation (DenseOoS) tailored for PanOoS. DenseOoS comprises a labeled test set of $1,000$ panoramic images with resolution $2048{\times}400$ and provides pixel-wise annotations for the PanOoS task, which greatly extend the semantic labels from DensePASS [ma2021densepass]. Relevant studies have shown that using generated outlier data for model training or evaluation is an effective approach [loiseau2025reliability, de2024placing]. Therefore, based on the DensePASS dataset, we use a state-of-the-art image generation technique, POC [de2024placing], to place multiple classes of outliers into images while generating annotation labels for these outliers to construct DenseOoS.
The normal labels cover $19$ classes that align with those of Cityscapes [cordts2016cityscapes], while the abnormal labels contain $30$ classes of outliers as demonstrated in Fig. 9. To ensure the high quality and plausibility of the generated images, particularly that the outliers are generated in realistic areas within the real street backgrounds, our screening process follows the “preliminary screening $\rightarrow$} independent screening $\rightarrow$} voting” workflow, eliminating unreasonable object locations, and normal situations. Specifically, a total of $14,031$ images are generated, of which $1,244$ are retained after preliminary manual screening. Following an independent screening and voting process by three examiners, we ultimately select $1,000$ images for the construction of the dataset. The outliers are diverse and distributed within contextually plausible regions of the images. Importantly, due to the $360^{\circ}$ panoramic layout, over $68\%$ of these outliers lie outside the typical $95^{\circ}$ narrow-FoV of pinhole cameras, which makes DenseOoS particularly suitable for the evaluation of the PanOoS task. More details and qualitative examples are provided in the Appendix.
@@ IV-C QuadOoS
To further explore the challenges of the PanOoS task in real-world applications, we collect panoramic images featuring outliers in unconstrained, unstructured, and dense-motion real scenes, and manually annotate these outliers to establish the first Quadruped Panoramic Out-of-distribution Segmentation dataset, QuadOoS. This dataset provides a crucial foundation for panoramic out-of-distribution segmentation research and facilitates the evaluation and optimization of PanOoS models in complex environments. We develop a mobile platform using a quadruped robot, and the robot is selected for its biomimetic gait, a complex form of locomotion found in nature, which introduces additional challenges for PanOoS. It can navigate vertical obstacles up to $15cm$ and inclines up to $30^{\circ}$, making it highly maneuverable in everyday environments. With $12$ joint motors, the robot replicates realistic walking motions at speeds up to $2.5m/s$. For sensing, we use a customized panoramic annular camera to capture a broader FoV ($360^{\circ}{\times}70^{\circ}$). The camera has a pixel size of $3.45{\mu}m{\times}3.45{\mu}m$, a resolution of $5$ million effective pixels. Mounted on the quadruped robot (see Fig. 10(a)), the camera ensures an optimal field of view.
@@
Using this platform, we conduct data collection in campus environments with outliers, creating the Quadruped Panoramic Out-of-distribution Segmentation (QuadOoS) dataset. In particular, we collect data across $8$ blocks in and around the campus, with each block averaging approximately $50m$ in length. The collected raw images consist of a total of $4,800$ frames ($600$ frames per block), each with an original size of $2048{\times}476$ pixels. To avoid misleading the model due to overexposed regions at the center of the lens and to maintain consistency with DenseOoS, we crop the upper portion of the original images (i.e., the sky), which does not affect fair evaluation. Subsequently, to ensure the diversity of outliers and the satisfactory imaging quality, we retain a final selection of $106$ original panoramic images after manual screening by five examiners, thereby constructing the QuadOoS dataset. Detailed information is provided in Fig. 10(b); see Appendix B-A for details of outlier types and examples from QuadOoS.
@@
Due to the gait of the quadruped robot, the collected panoramic images exhibit irregular shaking, particularly along the Y-axis (see Fig. 10(c)). Compared to panoramic images captured under static conditions, the images in QuadOoS present significant challenges to existing methods, such as uneven exposure, color inconsistencies due to the broader FoV, and increased motion blur, as rapid relative displacement between moving objects and the background intensifies the blurring effect. Each image contains an average of $7$ outliers, accurately reflecting real-world situations and posing substantial challenges for PanOoS methods. As illustrated in Table I, with a broader FoV ($360^{\circ}{\times}70^{\circ}$), QuadOoS significantly differs from traditional pinhole-OoS datasets [lis2019detecting, blum2021fishyscapes, hendrycks2019scaling, chan2021segmentmeifyoucan], and complements DenseOoS by enabling the exploration of motion-induced and panoramic-specific challenges in PanOoS.